OpenStudy (anonymous):

Evaluate. -3^2-5(-3)-8

OpenStudy (anonymous):

@dss08a

OpenStudy (abb0t):

9+15-8

OpenStudy (anonymous):

9+15-8 = 16

OpenStudy (anonymous):

when you square a negative it comes out a positive. but when you cube it, it stays a negative. just use PEMDAS for these type of questions. and screw you @abb0t

OpenStudy (accessdenied):

If this were (-3)^2 (noting parentheses), then it would become positive. (-3)^2 = 9 Because this has no parentheses, it would only apply to the 3 and thus: -3^2 = -9 It's akin to (-x)^2 vs. -x^2, (-x)^2 = (-1)^2 x^2 = x^2, but -x^2 the negative does not interact with the squared term (else y=-x^2 would look the same as y=x^2)

OpenStudy (accessdenied):

Thus we should obtain -9 + 15 - 8 = 15 - 17 = -2